1.A 4.50 hp water pump is pumping water from a well 7.50 m below ground. How many liters of water can it lift to the ground in 1 hour? 1 L of water has a mass of exactly 1 kg.
Answer #1

1 hp = 745.69 W

4.5 hp = 4.5*745.69 = 3355.649 W = 3355.649 J/s

1 hr = 1*60*60 = 3600 s

Energy Conservation

3355.649 J/s * 3600 s = mass* g * h

12080336.4 J = mass*9.81*7.5

mass = 12080336.4 / ( 9.81*7.5) = 164190.77 kg = 164190.77 L

liters of water can it lift to the ground in 1 hour = 164190.77 L
Answer #3

1 hp = 745.699872 watts

so

4.5 hp = 3355.64942 watts.

So work done by motor in 1 hour = Power * Time =  3355.64942*3600 = 12080337.912 joules

Work done to life water = m*g*h = work done by pump = 12080337.912 joules

m*9.81*7.5 = 12080337.912

m = 164190.797309 kg or litres ===>Answer
Answer #4

1 hp = 746 W

so 4.5 hp = 746*4.5 = 3357 W

1 watt means 1J/s

this means 1 joule of work is done in 1 seconds.

water is 7.5m below , so energy required to bring 1 kg(or 1L) to surface = mgh = 1*9.8*7.5= 73.5 J

power of pump is 3357 W , so in 1 second the water bring to ground by the pump = (3357/73.5) =45.67 kg

so in 1 hour (=3600s) = 45.67*3600 = 164412 kg

so in hour the litres of water pumped is 164412 L
Answer #5

First, lets calculate the work done in lifting 1l of water = 1kg from a depth of 7.5m

the energy is nothing but potential energy stored = mgh.

so word done in lifting 1l of water = 1*9.8*7.5 = 73.5 J

Now, the motor capacity is 4.5hp.

1 hp = 746J/s

So, in one second, the pump can do 746joules of work.

Which means, in one hour, the work done by the pump is

4.5*746*60*60 = 12085200 Joules.

So, dividing this by the work done to lift 1l of water, we have

12085200/73.5 = 164,424.5 litres of water in one hour.

assumption: g= 9.8 m/s2